OpenStudy (anonymous):

Can someone help me with understanding this 7√3 ?

OpenStudy (alexandervonhumboldt2):

this is a square root: 7sqrt3=sqrt(49*3)=\[\sqrt{147}\]

OpenStudy (alexandervonhumboldt2):

is there anything spesific you dont understand?

OpenStudy (anonymous):

Thank you! I am having to figure this up on a special right triangle in geometry, and I forgot what this meant.

OpenStudy (adi3):

any other questions

OpenStudy (anonymous):



OpenStudy (anonymous):

No. This is what I was having trouble with. I understand now (:

OpenStudy (adi3):

alright